OpenStudy (anonymous):

integration of sin ^{5}X

OpenStudy (anonymous):

you want answer in terms of cos3x, cos 5x or in terms of cos^3x , cos^5 x ?

OpenStudy (anonymous):

in terms of cos^3x , cos^5 x please:)

OpenStudy (anonymous):

well, if you want it in terms of cos^3 x and cos^5 x what you can do is, sin^5x = (sin^2 x)^2 . sin x and then write sin^2x as 1-cos^2x to get \(\large \int (1-\cos^2x)^2\sin x dx\) now you can put u= cos x du=.... ?

OpenStudy (anonymous):

yea good gohan :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

tell me if you have any doubts solving further.

OpenStudy (anonymous):

is there way of not using integration by substitution?

OpenStudy (anonymous):

if you can express sin^5 x in terms of first powers of sin 3x and sin 5x, then without substitution, you can get the answer, but in terms of cos 3x and cos 5x

OpenStudy (anonymous):

the answer is -cosx +2/3 cosx^3x-1/5cos^5x... but I don't know how the substitution works:/

OpenStudy (anonymous):

did you get what i explained ? did you try u=cos x ? du=..... ? if yes, where are u stuck ?

OpenStudy (anonymous):

is it du=-sinx dx?

OpenStudy (anonymous):

yes, so the integral turns to be \(\large -\int (1-u^2)du\) right ?

OpenStudy (anonymous):

*** \(\large -\int (1-u^2)^2du\)

OpenStudy (anonymous):

oh right!! thank you:))

OpenStudy (anonymous):

welcome ^_^